(Using Excel) In the Stanford Stadium problem with 60000 seats and two prices (General Public vs....

  1. (Using Excel) In the Stanford Stadium problem with 60000 seats and two prices (General Public vs. Student), suppose each General Public ticketholder also buys $20 in concessions resulting in a $10 contribution margin, and each Student ticketholder buys $10 in concessions resulting in a $5 contribution margin. What are the prices that maximize total contribution? Write out your formulation.

General Public: dg (p g) = (120000 - 3000 p g )+

Student:   ds (ps) = (20000 -1250ps) +
Answer #1

Marginal cost for general public, MCpg = $20

Marginal cost for students, MCps = $10

Demand of general public, dg = 120,000 - 3,000pg

pg = (120,000 - dg) / 3,000

Total revenue from general public, TRpg = pg x dg = (120,000dg - (dg)2) / 3,000

Marginal revenue from general public, MRpg = dTRpg / d(dg) = (120,000 - 2dg) / 3,000

Again,

Demand of students, ds = 20,000 - 1,250ps

ps = (20,000 - ds) / 1,250

Total revenue from students, TRpg = ps x ds = (20,000ds - (ds)2) / 1,250

Marginal revenue from students, MRps = dTRps / d(ds) = (20,000 - 2ds) / 1,250

Contribution margin is maximized when each group equates its MR with MC.

(i) MCpg = MRpg

20 = (120,000 - 2dg) / 3,000

60,000 = 120,000 - 2dg

dg = (120,000 - 60,000) / 2 = 30,000

pg = (120,000 - dg) / 3,000 = (120,000 - 30,000) / 3,000 = 90,000 / 3,000 = 30

(ii) MCps = MRds

10 = (20,000 - 2ds) / 1,250

12,500 = 20,000 - 2ds

ds = (20,000 - 12,500) / 2 = 3,750

ps = (20,000 - ds) / 1,250 = (20,000 - 3,750) / 1,250 = 16,250 / 1,250 = 13